a rectangle is 7 times as long as it is wide if the perimeter is 175 what are the dimensions
Answer by checkley77(12844)

L=7W
2L+2W=175
2*7W+2W+175
14W+2W=175
16W=175
W=175/16
W=10.9375 ANS.
L=7*10.9375=76.5625 ANS.
PROOF:
2*76.5625+2*10.9375=175
153.125+21.875=175
175=175